A water of tankis 1.2mm square and 1.35mm deep. If it is half full of water how many times can a 9 liter bucket be field from the tank

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

You might want to check your units
There is very little water with volume (1.2 mm)^2*1.35 mm=1.944 cc. Half full would be 0.97 cc or about 19 drops.
-
if it is 1.2 m square (1.44 m^2 area of the bottom)*1.35 m=1.944 m^3 of water
That is 1.944 m^3* 100cm/m*100cm/m*100 cm/m=1,944,000 cc or 1944 liters
Half of that is 972 liters so it would be 108 times filling the 9 l bucket.

Answer by ikleyn(52786)

.

Small note to the solution by  @boreal.

The volume of the  "tank"  is   1.2 mm * 1.2 mm * 1.35 mm = 1.944 mm^3  (cubic  MILLIMETERS,  not  cc  "cubic centimeters").


Feeling of numbers is of great help (!)